OpenStudy (angelina__nievse):

Can some one check my work I will medal and fan. Question 1 simplify (2^5/3^2)^4 my answer 2^9/3^8 Question 2 simplify 9^2/9^7 my answer is 1/9^5

OpenStudy (mathmate):

Q2 is correct: 9^2/9^7 = 9^(2-7)=9^(-5)=1/9^5 Q1 uses the law of exponents (a^x)^y=a^(xy) So \((2^5/3^2)^4=(2^{5*4}/(3^{2*4}) =\)... I'll let you complete it. Your denominator was correct.

OpenStudy (angelina__nievse):

Would it be 2^20/3^8

OpenStudy (mathmate):

Exactly, well-done! :)